### v2

# Schrödinger’s principle eliminates the EPR-locality paradox

Abstract
We introduce a principle, implicitly contained in Schrödinger’s paper [Schr35], which allows a proof of the non-existence of the EPR-locality paradox in the Copenhagen interpretation of quantum mechanics. The paradox is shown to be well-posed already in the case of the simplest example of an entangled state of two spins one-half. We assume that any measurement results in the collapse of the wave-packet.
1 Introduction
The present year marks the centennary of (Schrödinger) quantum mechanics, which originates on his papers in the journal Annalen der Physik (the first one being Ann. Phys. 79, 361 (1926)). It seems therefore natural to inquire today on the axioms of quantum mechanics, in the words of Wigner [Wi60] “as formulated explicitly by the great physicist Dirac. There are two basic concepts in quantum mechanics: states and observables. The states are vectors in Hilbert space, the observables self-adjoint operators acting on these vectors.” We may ask whether any doubts remain about the general validity of this framework: for instance, is the EPR-locality paradox [EPR35] cleared up?
In the present paper we attempt to complete some steps in Schrödinger’s analysis [Schr35] which allow a complete proof of the non-existence of the EPR paradox in the case of measurements of nonlocal quantities, such as spin components or photon polarizations, within the Copenhagen interpretation. Griffiths [Gri11] has arrived at the same conclusion within his own probabilistic interpretation, see also [Gri19]. The issue of local measurements is a different one, see Haag’s book ([Ha22], p.107 and p.316), and the work of Fewster and Verch [FV25]. Our objective is more modest: we wish to emphasize that the problem was already solved by Schrödinger, but the lack of the missing steps within even the simplest context may give the general false impression that his approach was vague or not sufficiently precise.
Although Haag’s conclusion, formulated in terms of probabilities, (see, in particular, [Ha22], p. 107), is equivalent to ours, the simple Wigner-Dirac approach mentioned above suggests that the main issues occur already in the case of quantum systems with finite number of degrees of freedom (ndof), for which there are only two guiding principles which rule the structure of wave-vectors: the building of tensor products and the superposition principle. The principle which we suggest to call Schrödinger’s principle concerns the formation of tensor products.
In fact, as we shall see, the (so-called) EPR-locality paradox is well-posed already in the simplest example of states of two spins one-half. In general, for a finite region $\Lambda\subset\mathbf{Z}^{\nu}$, where $\nu$ denotes the dimension, the Hilbert space is given by ${\cal H}_{\Lambda}=\otimes_{x\in\Lambda}\mathbf{C}_{x}^{2}$, and the set of observables is the set ${\cal A}(\Lambda)=B({\cal H}_{\Lambda})$ of bounded operators on this Hilbert space, generated by the set of operators $S^{i}_{x}=1/2\sigma^{i}_{x}$, where $i=1,2,3$ and $\sigma^{i}_{x}$ denote the Pauli matrices at the site $x$.
2 Generalities on states and observables
Let $\rho_{\Lambda}$ denote a density matrix on ${\cal H}_{\Lambda}$, that is, a positive, Hermitean matrix of unit trace. A state $\omega_{\Lambda}$ on ${\cal A}(\Lambda)$ is a positive, normed linear functional on ${\cal A}(\Lambda)$: $\omega_{\Lambda}(A)=Tr_{{\cal H}_{\Lambda}}(\rho_{\Lambda}A)\mbox{ for }A\in{\cal A}(\Lambda)$ (positive means $\omega_{\Lambda}(A^{{\dagger}}A)\geq 0$, normed $\omega_{\Lambda}(\mathbf{1})=1$.).
A state $\omega_{\Lambda}$ on the observable algebra ${\cal A}_{\Lambda_{1}}\cup{\cal A}_{\Lambda_{2}}$ is a product state iff there are states $\omega_{\Lambda_{1}}$ on ${\cal A}_{\Lambda_{1}}$ and $\omega_{\Lambda_{2}}$ on ${\cal A}_{\Lambda_{2}}$ such that
$\omega_{\Lambda}(XY)=\omega_{\Lambda_{1}}(X)\omega_{\Lambda_{2}}(Y)\forall X\subset{\Lambda_{1}}\mbox{ and }Y\subset\Lambda_{2}$
A state is called separable iff it is a convex combination of product states. Otherwise it is called entangled. The former states are thereby “classical” in the sense of exhibiting “product correlations”, and the entangled states, in the complementary set, are supposedly the most “intrinsically quantum”. Notable examples of entangled states called Bell states [Li14], one of the most famous of them being
$\omega_{B}(.)\equiv(\Psi_{B,-},.\Psi_{B,-})\mbox{ where }\Psi_{B-}\in\mathbf{C}_{1}^{2}\otimes\mathbf{C}_{2}^{2}$ (1)
where
$\Psi_{B,-}\equiv\frac{1}{\sqrt{(}2)}(|+)_{1}\otimes|-)_{2}-|-)_{1}\otimes|+)_{2})$ (2)
and $|\pm)_{1,2}$ denote the basis of the two orthonormal eigenvectors of $\sigma^{3}$ on the spaces $\mathbf{C}_{1}^{2}$ resp. $\mathbf{C}_{2}^{2}$.
Quantum spin systems are similar to quantum fields because of the now famous Lieb-Robinson bound [LR72]: for $A,B\in{\cal A}_{0}$, where ${\cal A}_{0}$ denotes the algebra of observables at the origin,
[…]
$|x|<v_{g}|t|$ (4)
where the quantity $v_{g}$ is called the group velocity, which satisfies
$0<v_{g}<\infty$ (5)
for a large class of interactions, which include the finite-range interactions (see [BRo297], p. 254).
3 The EPR-locality paradox and Schrödinger’s principle
We refer to the observers associated to the “soap-opera” ([Li14], p.34) as Alice, who performs measurements of $\sigma^{3}_{2}$ and Bob, who measures $\sigma^{3}_{1}$. In addition, Charlie, a third character in the opera, introduced by Griffiths ([Gri19], p.19), prepares a spin singlet at the center of the laboratory and sends the two particles in opposite directions towards two apparatuses 1 and 2, constructed, respectively, by Alice and Bob, who are both aware of the initial state  (2). This spin system has no dynamics, but its observables may be considered as part of an infinite variety of (infinite) quantum spin system, with a definite dynamics, see, e.g., [Wre22]. Take any one of them as an example, the results are all, of course, equivalent, because they depend only on the validity of the Lieb-Robinson bound  (3). The corresponding symmetries of time translation and space translation are, then, described by so-called automorphisms $\tau_{t}$ and $\tau_{x}$, respctively, of an algebra $\cal A$} of the infinite system, which are one-to-one mappings of $\cal A$} onto itself, which preserve the algebraic structure. Above, $t\in\mathbf{R}$ and $x\in\mathbf{Z}^{\nu}$, where $\nu$ denotes the dimension.
We now ask: in what consists the EPR-locality paradox [EPR35], and is it well posed? The paradox consists in the following consideration: if Alice performs a measurement and finds a certain value of the (3-component) of the spin, she is able to deduce the value that Bob will be measuring, even if he is beyond the reach of any signal, due to the fact that she knows the initial state ( (1) or  (2)).
Notice that in  (2), the two basic building blocks of quantum mechanics - tensor products and the superposition principle- are present: we refer to Lieb’s [Li14] stimulating lectures on the subject, which we follow in part. About this state, Schrödinger [Schr35] writes: “by their interaction, the two representatives (or functions) have become entangled.”. He says, in addition, that “entanglement is not one, but the characteristic trait of quantum mechanics, the one that enforces its entire line of departure from classical lines of thought”.
We formulate Schrödinger’s description of this “trait” as an independent principle:
Schrödinger’s principle: The characteristic “trait” of quantum mechanics is that two physical systems that are brought into temporary interaction and then completely separated, can no longer be described, even after the interaction has utterly ceased, by individual wave-functions for each of the two systems.
The word “interaction” used by Schrödinger in Schrödinger’s principle refers to the formation of an entangled state such as  (1), the “complete separation” is the operation performed by Charlie. We have, now:
Proposition 3.1.
The EPR-locality paradox is well posed under the following assumptions: 1) the time $T$ of a quantum mechanical measurement satisfies $T<\infty$ ([Wre23]) If, in addition, we assume that 2) at the moment the measurement is performed, there occurs the collapse of the wave-packet then the EPR-locality paradox disappears as soon as Schrödinger’s principle is invoked.
Proof.
Under assumption 1), if the initial state is subject to a dynamics satisfying the assumptions of the Lieb-Robinson bound  (3), there is a finite group velocity FGV  (4), and therefore Bob’s signal cannot reach Alice if he is at a distance greater than $v_{g}T$ from Alice. The paradox is thus well-posed, the contradiction been now with FGV, not with relativity.
We now come to part 2.). Schrödinger’s principle implies that any measurement on Bob’s spin or Alice’s spin is to be regarded as a measurement on the entire two-spin system. When Alice finds that her spin points “up”, there has been, by assumption 2), a collapse of the state $\Psi_{B,-}$ to the state $|-)_{1}\otimes|+)_{2}$, whereby she actually learns that the two-spin system is in the state $|-)_{1}\otimes|+)_{2}$.
Therefore, she does learn at the same time that Bob will be necessarily measuring the value -1/2. This result strongly contrasts with EPR’s version of the experiment, which may be described as follows: at the moment of her measurement, Alice is able to deduce from her knowledge of the initial state  (2) that Bob will be measuring the value 1/2-1/2, thereby obtaining a new information, which, as we have previously shown, violates FGV if Bob is sufficiently far away.
4 Conclusion
As remarked in the introduction, our contribution in this paper was only to show that the solution in Schrödinger’s paper [Schr35] of the EPR-locality paradox, within the Copenhagen interpretation, may be made completely precise. Griffiths’ solution ([Gri11], [Gri19]) was obtained within his probabilistic approach, but Griffiths does not make explicit use of the collapse of the wave-packet, as he himself remarks in [Gri19].
Acknowledgement We should like to thank Elliott Lieb very heartily. He has been a mentor and a friend for several decades, and his stimulating lectures [Li14] have been the point of departure of the present manuscript. Thanks also to Klaas Landsman for constructive criticism of a previous version, and for reminding us of the important work of Haag on the subject, as well as of Fewster and Verch.